# Gas supply shocks, Uncertainty and Price setting: Evidences from Italian Firms

### v3

@@ 4 Conclusions
This paper provides new evidence on how natural gas supply shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. We identify a gas supply shock through an external IV-VAR framework and we analyze its implications for firms’ pricing using quarterly survey data from the Bank of Italy’s SIGE, spanning over two decades. Our analysis reveals three main findings. First, natural gas supply shocks emerge as a significant driver of electricity prices in Italy. An 1% increase in the spot natural gas price measured by the Dutch TTF future determines a 0.5% increase in the Italian electricity price, measured by the PUN on impact. Second, firms respond to natural gas supply innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock, with effects dissipating within 5-6 quarters, consistent with the transitory nature of energy price shocks documented in the literature. The increase of inflation uncertainty in response to the shock suggests that energy price disruptions affect not only the level of prices but also firms’ ability in forecasting future inflation. Third, and most importantly, we document substantial non-nonlinearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weak demand conditions prevent cost pass-through.

### v4

@@ 4 Conclusions
This paper provides new evidence on how natural gas supply shocks affect firms’ pricing decisions and inflation expectations in Italy, a country heavily reliant on natural gas as its primary energy source. We identify a gas supply shock through an external IV-VAR framework and we analyze its implications for firms’ pricing using quarterly survey data from the Bank of Italy’s SIGE, spanning over two decades. Our analysis reveals three main findings. First, natural gas supply shocks emerge as a significant driver of electricity prices in Italy. An 1% increase in the spot natural gas price measured by the Dutch TTF future determines a 0.5% increase in the Italian electricity price, measured by the PUN. Second, firms respond to natural gas supply innovations by adjusting both their current and expected own prices, with inflation uncertainty rising following the shock. These effects take about one year to dissipate fully, consistently with the transitory nature of energy price shocks documented in the literature. Hence, the increase of inflation uncertainty in response to the shock suggests that energy price disruptions affect not only the level of prices but also firms’ ability in forecasting future inflation. Additionally, firms respond to the shock adjusting their prices more aggressively. Additionally, substantial differences emerge across sectors, with adjusting their prices more aggressively with respect to services sector firms. Third, we document substantial nonlinearities in firms’ responses that depend on the pre-existing level of inflation uncertainty. When uncertainty is high, firms successfully pass through cost increases to consumers, maintaining elevated prices throughout the adjustment period. In contrast, when uncertainty is low, the recessionary effects of the shock dominate, causing firms to reduce prices below baseline as weakened demand conditions prevent cost pass-through.
[…]
Appendix A Uncertainty index
Figure A.1 plots the inflation-uncertainty index, constructed following binder2017measuring, separately for three groups of firms: (i) those that were informed about current inflation at the time of the interview; (ii) those that received no information about current inflation; (iii) those that were informed only about Euro Area inflation.
Because SIGE started interviewing firms without providing current-inflation information only in the later part of the sample, the uncertainty series for groups (ii) and (iii) are substantially shorter. Nonetheless, uncertainty in these groups exhibits broadly similar dynamics to group (i). The share of firms classified as uncertain in case (ii) is about five percentage points higher on average, but this gap narrows sharply during the post-pandemic recovery. This pattern indicates that the surge in inflation uncertainty in that period is not primarily driven by informational frictions.
Figure A.1: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring. Vertical red bars highlight moments of high uncertainty. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
Appendix B IV-VAR robustness
B.1 Controlling for geopolitical risk
Figure B.2 shows the impulse response functions obtained by augmenting our energy-sector VAR with the geopolitical risk index of caldara2022measuring and re-estimating the model using the same external instrument. The responses are virtually unchanged relative to the baseline specification. This result indicates that the instrument is not simply capturing broad geopolitical risk shocks and supports its interpretation as identifying disruptions to the physical supply of natural gas.
Gas supply shock
[Uncaptioned image] Figure B.2: Impulse response functions of energy market variables to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019, caldara2022measuring. Sample: 2004:M1–2025:M6.
Appendix C BVAR robustness
C.1 Removing 2020 from the sample
Gas supply shock
[Uncaptioned image] Figure C.3: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Responses are computed by removing observations corresponding to year 2020 from the sample. Shaded areas represent 68% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
C.2 Jurado measure of firms’ inflation uncertainty
Gas supply shock
[Uncaptioned image] Figure C.4: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, computed with the methodology of Jurado2015, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations, barigozzi2024large. Sample: 2005:Q1–2025:Q2.